# XLD: A Cross-Lane Dataset for Benchmarking Novel Driving View Synthesis

### v2

@@ Abstract
Thoroughly testing autonomy systems is crucial in the pursuit of safe autonomous driving vehicles. It necessitates creating safety-critical scenarios that go beyond what can be safely collected from real-world data, as many of these scenarios occur infrequently on public roads. However, the evaluation of most existing NVS methods relies on sporadic sampling of image frames from the training data, comparing the rendered images with ground truth images using metrics. Unfortunately, this evaluation protocol falls short of meeting the actual requirements in closed-loop simulations. Specifically, the true application demands the capability to render novel views that extend beyond the original trajectory (such as cross-lane views), which are challenging to capture in the real world. To address this, this paper presents a novel driving view synthesis dataset and benchmark specifically designed for autonomous driving simulations. This dataset is unique as it includes testing images captured by deviating from the training trajectory by $1-4$ meters. It comprises six sequences encompassing various time and weather conditions. Each sequence contains $450$ training images, $120$ testing images, and their corresponding camera poses and intrinsic parameters. Leveraging this novel dataset, we establish the first realistic benchmark for evaluating existing NVS approaches under front-only and multi-camera settings. The experimental findings underscore the significant gap that exists in current approaches, revealing their inadequate ability to fulfill the demanding prerequisites of cross-lane or closed-loop simulation. Our dataset is released publicly at the project page: https://3d-aigc.github.io/XLD/.
@@ 1 Introduction
Autonomous driving (AD) simulation, which bridges the gap between the real world and the virtual world, is essential for the testing and development of on-vehicle autonomous driving software [7]. Research indicates that employing effective simulation methods can significantly expedite the evaluation of safety tests for autonomous driving, achieving a speedup of approximately $10^{3}$ to $10^{5}$ times faster compared to real-world testing [20]. This compelling evidence underscores the importance of leveraging simulation to enhance the efficiency and effectiveness of autonomous driving development. However, the self-driving industry primarily conducts system testing using two approaches: log replay, which involves testing on pre-recorded real-world sensor data, and real-world driving, where new miles are driven to gather additional data for testing purposes [58].
@@
In closed-loop simulation, it is essential for the vehicle to have the freedom to respond to control commands within the simulation environment, rather than being confined to strictly following the original trajectory from logs. To corporate with the recently developed end-to-end autonomous driving [61, 23], designing a neural simulator for AD simulation [58, 54, 52], which can render photo-realistic images on novel views for closed-loop simulation and algorithm training, is in high demands. The main scientific problems boil down to the 3D reconstruction [33] and novel view synthesis (NVS) [56, 54], which are also long-standing problems in computer vision and computer graphics. Traditional methods such as [40, 41] have dominated the major deployment of 3D scene reconstruction for a long time. However, those reconstructed scenes cannot be directly used to produce photo-realistic novel views, thus imposing large restrictions on sensor simulations. As a result, the industrial bridges the sim-to-real gap by parametric and procedural modeling technique [47] or human-involved creations. With the recent rapid development of 3D implicit field, such as neural radiance field (NeRF [35]) and explicit primitive representations, i.e., 3D Gaussian Splatting (3DGS [25]), reconstructing a scene from a collection of images serves as the foundation of end-to-end autonomous driving simulation [22]. These techniques enable rendering high-quality and photorealistic images on novel views.
@@
Presently, the majority of approaches evaluate the performance of novel view synthesis (NVS) results by splitting the dataset into training and testing sets. However, this strategy of splitting and sampling leads to an interpolation benchmark, which we argue is insufficient for evaluating whether the trained models can effectively render simulation-ready (i.e., cross-lane) and high-fidelity data for closed-loop simulations. On the contrary, our proposed DBN dataset is a brand-new benchmark that evaluates the synthesis quality on a cross-lane view with additionally captured GT images. Our dataset and benchmark focus on assessing the NVS capability specifically for cameras in cross-lane scenes. The primary objective is to evaluate the performance of cameras in generating accurate and realistic novel views in scenarios involving multiple lanes. Specifically, we introduce the XLD dataset, which encompasses the generation of $150\times 3$ rendered images for each scene. Additionally, we evaluate cross-lane novel view synthesis by rendering $50$ images with deviations of $1$m, $2$m, and $4$m from the training trajectory. Using the XLD dataset, we conduct a benchmark of leading methods, which are based on either NeRF or 3DGS, using well-established NVS metrics. The results of our benchmark demonstrate that our proposed dataset offers a comprehensive evaluation benchmark tailored specifically to the requirements of closed-loop simulation. Furthermore, our benchmarking results reveal intriguing findings, emphasizing the value of the proposed dataset.
@@ 2.1 Autonomous Driving Simulation
In the past few years, there has been a surge in the use of autonomous driving simulations [29]. These simulators are instrumental in validating planning and control mechanisms, producing educational and evaluative datasets, and significantly cutting down the time needed to perform these functions. The current landscape is dominated by two predominant categories of simulation tools: model-based and data-driven. Model-based simulation platforms, such as PyBullet [18] MuJoCo [46], AirSim [42] and CARLA [19], utilize advanced computer graphics to replicate vehicles and their surroundings. However, the manual effort required to construct these models and to program the dynamics of the vehicles can be quite demanding and lengthy. Moreover, the visual output may sometimes fall short of the necessary realism, which can adversely affect the efficacy of perception systems when they are put into operation.
Previously, NVS heavily relied on conventional image processing techniques. For instance, Chaurasia et al.[12] propose the use of depth synthesis from over-segmented graph structures, while AADS[28] employs filtered and completed dense depth maps for warping novel view images through image stitching. A data-driven simulation platform VISTA [3, 2] leverages datasets from the real world to create comprehensively labeled and hyper-realistic simulations. In more recent times, there has been a wave of innovations that employ the NeRF method to superficially simulate driving perspectives. These new approaches excel in the creation of hyper-realistic images and have been shown to surpass traditional view synthesis algorithms in the realm of autonomous driving simulation. Recent advances in neural novel view synthesis significantly accelerate the rapid development of the next-gen driving simulation, which exhibits superior expressiveness and flexibility compared to traditional methods. Our dataset and benchmark work is specifically designed for those methods.
@@ 2.2 NeRF-based NVS for Driving Simulation
Motivated by the NeRF-based methods and point-based differentiable rendering [59, 1, 39, 30], 3D Gaussian Splatting (3DGS) [25] opens a new era with the leading advantages in explicit representation and real-time rendering capability. Within a concise timeframe, numerous methods [14, 56] have emerged that focus on road scene reconstruction and NVS by leveraging the 3DGS representation. For instance, PVG [14, 27] introduces periodic vibration-based temporal dynamics to reconstruct dynamic urban scenes. StreetGaussian [56] models the dynamic urban street environment as a collection of point clouds with semantic logits and 3D Gaussians, each associated with either a foreground vehicle or the background. DrivingGaussian [62] uses incremental static 3D Gaussians to represent the scene’s static background. It also employs a composite dynamic Gaussian graph to handle multiple moving objects with LiDAR data.  [31] introduce a hybrid method that combines radiance fields with 3DGS representation, effectively eliminating the requirement for point initialization in the context of urban scene NVS.  [53] propose a hybrid Gaussian representation specifically designed for performing online dense mapping in unbounded large-scale scenes. GaussianPro [15] utilizes priors from reconstructed scene geometries and patch-matching techniques to generate precise Gaussians, leveraging the scene’s existing structure. DC-Gaussian [50] introduces adaptive image decomposition for modeling reflections and occlusions. It incorporates illumination-aware obstruction modeling to handle reflections and occlusions under varying lighting conditions in urban scene novel view synthesis. Our dataset and benchmark specifically focus on evaluating the performance of neural-based driving simulation in NVS, particularly in cross-lane scenarios. A portion of the mentioned methods serve as our baselines, taking into account the code availability. Moreover, generalizable 3D-GS methods such as PixelSplat [11], Mvsplat [13], and GGRt [26] also attempts to synthesis novel images within a well-pretrained generalizable feed-forward Gaussian networks.
Figure 1: Our datasets encompass six distinct scenes, where each scene involves the vehicle following an on-road trajectory. To generate training data for the cameras and LiDAR sensor, we sample $150$ waypoints along each trajectory. The trajectory is visually emphasized using the color red.
@@ 2.4 Datasets in Autonomous Driving
In autonomous driving training and benchmarking, there exist a number of datasets available. For example, KITTI [21]. KITTI-360 [32], vKITTI [8], CityScapes [17], Mapillary [37], ApolloScape [24], Waymo Open Dataset [43]. nuScenes [9], Argoverse [10] and Argoverse 2 [51]. BDD100K [60]. OpenLane-V2 [49], etc.. Those previous works have laid the groundwork for research and development in autonomous driving algorithms. A comprehensive survey of datasets related to autonomous driving refers to [34]. However, as far as our knowledge goes, none of the existing research has specifically addressed the evaluation of novel view synthesis techniques tailored for autonomous driving simulation, particularly in terms of their ability to meet the high demands of cross-lane capability.
@@ 3 Dataset
Unlike previous datasets (both real-world and synthetic) that only capture single trajectories on the road, our dataset also captures several trajectories parallel to the main trajectories to better evaluate the novel-view render quality. To generate cross-lane data in the created worlds, we utilize Carla [19], a autonomous driving simulator platform built on Unreal Engine.
@@ 3.1 Sensor Setup
To meet the needs of most NeRF and 3D-GS algorithms, the sensors used to capture data include three color cameras (i.e. ‘left-front’, ‘front’, ‘right-front’) and one 3D laser scanner. The spatial relationships between sensors and the vehicle are fixed as shown in Fig. 2.
All three RGB cameras share identical intrinsic, lens parameters. Specifically, they have a resolution of $1920\times 1280$ pixels, a field of view (FOV) of $49.5^{\circ}\times 36.7^{\circ}$, a sensitivity of ISO 100, and a shutter speed of 5ms. The 3D laser scanner features 64 laser beams, a scanning range of 60 meters, and laser beam angles ranging from $-30^{\circ}$ to $30^{\circ}$ in vertical direction.
@@ 3.2 Data Generation
Our simulation environment consists of six scenes ("Town01," "Town02," "Town03," "Town04," "Town05," and "Town10", all of them are under CC-BY License provided by [19]) with various weather conditions, such as sunshine and rain, closely resembling real-world settings. An overview of all the scenes and the trajectories is shown in Fig. 1. Take one scene for example, the training set samples 150 times to capture three cameras’ images, LiDAR points, and the vehicle’s extrinsic information. The vehicle forward distance between two samples is 2 about meters. The test set includes four groups captured by the ’front’ camera. These images are aligned parallel to the training set, with each group exhibiting a progressive deviation of 0m, 1m, 2m, and 4m along the y-axis in vehicle coordinate, as shown in Fig. 2. Additionally, a few novel-view-synthesis for AD methods like GaussianPro [15], MARS [54], and UC-NeRF [16] need annotated sky masks to split the scene into foreground-sky and model the color compensation separately. We employ a pre-trained SegFormer [55] to effectively infer semantic segmentation masks and extract sky masks from them.
Figure 2: The composition of our training set and testing set. The training set consists of three RGB cameras (’front’, ’left-front’, ’right-front’) mounted on our vehicle lane. For the test set, we sample image sequences along trajectories that run parallel to the vehicle’s route. This encompasses four distinct test trajectories, each offset from the vehicle’s trajectory by 0 meters, 1 meter, 2 meters, and 4 meters, respectively. The sampling interval for the test set is five times the sampling interval used for the training set.
@@ 4 Benchmark
4.1 Benchmarking Environment
In order to thoroughly assess the performance and computational efficiency of the evaluating methods, we carried out a series of experiments utilizing an NVIDIA Tesla V100 16GB GPU. Our method was extensively validated across five distinct datasets featuring driving scenarios. We have presented our findings through both qualitative and quantitative assessments, which underscore the excellent performance and efficiency when juxtaposed with alternative methodologies.
4.2 Benchmarking Methods
MARS [54]: We inherit most of the parameter settings as the original MARS implementation. We employ the RAdam optimizer with $1\times 10^{-3}$ learning rate. Due to the fact that our scenes are static without moving objects, we disable $\mathcal{L}_{sem}$, the rest of loss functions remain the same. We train the model with 50,000 steps.
UC-NeRF [16]: We inherit most of the parameter settings as the original UC-NeRF implementation. We employ the AdamW optimizer with $2.5\times 10^{-3}$ learning rate. The weight of sky loss is set to $2\times 10^{-3}$. We train the model with 40,000 steps.
3DGS [25]: We use the implementation of NerfStudio to evaluate our dataset. We employ the AdamW optimizer with $1\times 10^{-3}$ learning rate. For stability, we “warm-up” the computation in lower resolution. Specifically, we start the optimization using 4 times smaller image resolution and we upsample twice after 500 and 1000 iterations. We train the model with 30,000 steps.
PVG [14]: We employ the Adam optimizer and maintain a similar learning rate for most parameters as the original PVG implementation while we adjust the learning rate of the amplitude $\mathbf{A}$, opacity decaying $\mathbf{\beta}$ and opacity $\mathbf{O}$ to $3\times 10^{-5}$, $0.02$ and $0.005$ respectively. We train the model with 30,000 steps.
GaussianPro [15]: In alignment with the approach described in GaussianPro, our models are trained for 30,000 iterations across all scenes following GaussianPro’s training schedule and hyperparameters. The interval step of the progressive propagation strategy is set to 50 where propagation is performed 3 times. The threshold $\sigma$ of the absolute relative difference is set to 0.8. For the planar loss, we set $\beta=0.001$ and $\gamma=0.001$.
4.3 Used Metrics
We adopt the evaluation criteria employed by the aforementioned methods, which comprise Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) as our evaluation metrics. Furthermore, we ensure transparency and clarity by providing a detailed description of the experimental framework employed for comparing the methods.
Table 1: Results on our proposed dataset with the different offsets using front-only camera.

### v3

@@ Abstract
Comprehensive testing of autonomous systems through simulation is essential to ensure the safety of autonomous driving vehicles. This requires the generation of safety-critical scenarios that extend beyond the limitations of real-world data collection, as many of these scenarios are rare or rarely encountered on public roads. However, evaluating most existing novel view synthesis (NVS) methods relies on sporadic sampling of image frames from the training data, comparing the rendered images with ground-truth images. Unfortunately, this evaluation protocol falls short of meeting the actual requirements in closed-loop simulations. Specifically, the true application demands the capability to render novel views that extend beyond the original trajectory (such as cross-lane views), which are challenging to capture in the real world. To address this, this paper presents a synthetic dataset for novel driving view synthesis evaluation, which is specifically designed for autonomous driving simulations. This unique dataset includes testing images captured by deviating from the training trajectory by $1-4$ meters. It comprises six sequences that cover various times and weather conditions. Each sequence contains $450$ training images, $120$ testing images, and their corresponding camera poses and intrinsic parameters. Leveraging this novel dataset, we establish the first realistic benchmark for evaluating existing NVS approaches under front-only and multicamera settings. The experimental findings underscore the significant gap in current approaches, revealing their inadequate ability to fulfill the demanding prerequisites of cross-lane or closed-loop simulation. Our dataset and code are released publicly on the project page: https://3d-aigc.github.io/XLD.
@@ 1 Introduction
Autonomous driving (AD) simulation, which bridges the gap between the real and virtual worlds, is essential for testing and developing autonomous driving software in vehicles [7]. Research indicates that employing effective simulation methods can significantly expedite the evaluation of safety tests for autonomous driving, achieving a speedup of approximately $10^{3}$ to $10^{5}$ times faster than real-world testing [21]. This compelling evidence underscores the importance of leveraging simulation to enhance the efficiency and effectiveness of autonomous driving development. However, the self-driving industry primarily conducts system testing using two approaches: log replay, which involves testing on pre-recorded real-world sensor data, and real-world driving, where new miles are driven to gather additional data for testing purposes [61]. In closed-loop simulation, the vehicle must be free to respond to control commands within the simulation environment rather than strictly following the original trajectory from logs. To promote the rapid advancement of end-to-end autonomous driving systems [64, 24], designing a neural simulator for AD simulation [61, 57, 55], which can render photo-realistic images on novel views for closed-loop simulation and algorithm training, is in high demands. The main scientific problems boil down to the 3D reconstruction [35], and novel view synthesis (NVS) [59, 57], which are also long-standing problems in computer vision and computer graphics. Traditional methods such as [43, 44] have dominated the major deployment of 3D scene reconstruction for a long time. However, those reconstructed scenes cannot be directly used to produce photo-realistic novel views, thus imposing large restrictions on sensor simulations. As a result, the industrial bridges the sim-to-real gap by parametric and procedural modeling technique [50] or human-involved creations. With the recent rapid development of 3D implicit fields, such as neural radiance field (NeRF [37]) and explicit primitive representations, i.e., 3D Gaussian Splatting (3DGS [26]), reconstructing a scene from a collection of images serves as the foundation of end-to-end autonomous driving simulation [23]. These techniques enable the rendering of high-quality and photorealistic images on novel views.
@@
Presently, most approaches evaluate the performance of NVS results by splitting the dataset into training and testing sets. However, this strategy of splitting and sampling leads to an interpolation benchmark, which we argue is insufficient for evaluating whether the trained models can effectively render simulation-ready (i.e., cross-lane) and high-fidelity data for closed-loop simulations. On the contrary, our proposed XLD dataset is a brand-new benchmark that evaluates the synthesis quality on a cross-lane view with additionally captured GT images. Our dataset and benchmark focus on assessing the NVS capability specifically for cameras in cross-lane scenes. The primary objective is to evaluate the performance of cameras in generating accurate and realistic novel views in scenarios involving multiple lanes. Specifically, we introduce the XLD dataset, which encompasses the generation of $150\times 3$ rendered images for each scene. Furthermore, we evaluate novel cross-lane view synthesis by rendering $30$ images with deviations of 00m, $1$m, $2$m, and $4$m from the training trajectory. Using the XLD dataset, we conduct a benchmark of leading methods, which are based on either NeRF or 3DGS, using well-established NVS metrics. Our benchmark results demonstrate that the proposed dataset offers a comprehensive evaluation benchmark tailored specifically to the requirements of closed-loop simulation. Furthermore, our benchmarking results reveal intriguing findings, emphasizing the value of the proposed dataset.
@@ 2.1 Autonomous Driving Simulation
In the past few years, there has been a surge in the use of autonomous driving simulations [31]. These simulators are instrumental in validating planning and control mechanisms, producing educational and evaluative datasets, and significantly cutting down the time needed to perform these functions. The current landscape is dominated by two predominant categories of simulation tools: model-based and data-driven. Model-based simulation platforms, such as PyBullet [18] MuJoCo [49], AirSim [45] and CARLA [19], utilize advanced computer graphics to replicate vehicles and their surroundings. However, the manual effort required to construct these models and program the vehicles’ dynamics can be quite demanding and lengthy. Moreover, the visual output may sometimes fall short of the necessary realism, which can adversely affect the efficacy of perception systems when they are put into operation.
Previously, NVS heavily relied on conventional image processing techniques. For instance, Chaurasia et al.[12] propose using depth synthesis from over-segmented graph structures. At the same time, AADS[30] employs filtered and completed dense depth maps for warping novel view images through image stitching. A data-driven simulation platform VISTA [3, 2] leverages datasets from the real world to create comprehensively labeled and photorealistic simulations. Recently, a wave of innovations has employed the NeRF method to simulate driving perspectives superficially. These new approaches excel in creating photorealistic images and have been shown to surpass traditional view synthesis algorithms in the realm of autonomous driving simulation. Recent advances in neural novel view synthesis significantly accelerate the rapid development of the next-gen driving simulation, which exhibits superior expressiveness and flexibility compared to traditional methods. Our dataset and benchmark are specifically designed for those methods.
@@ 2.2 NeRF-based NVS for Driving Simulation
Figure 1: Our datasets encompass six distinct scenes, each involving the vehicle following an on-road trajectory. To generate training data for the cameras and LiDAR sensor, we sample $150$ waypoints along each trajectory. The trajectory is visually emphasized using the color red.
@@
Motivated by the NeRF-based methods and point-based differentiable rendering [62, 1, 42, 32], 3D Gaussian Splatting (3DGS) [26] opens a new era with the leading advantages in explicit representation and real-time rendering capability. Within a concise timeframe, numerous methods [13, 59] have emerged that focus on road scene reconstruction and NVS by leveraging the 3DGS representation. For instance, PVG [13, 29] introduces periodic vibration-based temporal dynamics to reconstruct dynamic urban scenes. StreetGaussian [59] models the dynamic urban street environment as a collection of point clouds with semantic logits and 3D Gaussians, each associated with either a foreground vehicle or the background. DrivingGaussian [65] uses incremental static 3D Gaussians to represent the scene’s static background. It also employs a composite dynamic Gaussian graph to handle multiple moving objects with LiDAR data. HO-Gaussian [33] introduces a hybrid method that combines radiance fields with 3DGS representation, eliminating the requirement for point initialization in urban scene NVS. HGS-Mapping [56] proposes a hybrid Gaussian representation specifically designed for performing online dense mapping in unbounded large-scale scenes. GaussianPro [16] utilizes priors from reconstructed scene geometries and patch-matching techniques to generate precise Gaussians, leveraging the scene’s existing structure. DC-Gaussian [53] introduces adaptive image decomposition for modeling reflections and occlusions. It incorporates illumination-aware obstruction modeling to handle reflections and occlusions under varying lighting conditions in urban scene novel view synthesis. Our dataset and benchmark specifically focus on evaluating the performance of neural-based driving simulation in NVS, particularly in cross-lane scenarios. A portion of the mentioned methods serve as our baselines, taking into account the code availability. Moreover, generalizable 3D-GS methods such as PixelSplat [11], Mvsplat [14], and GGRt [28] also attempts to synthesize novel images within a well-trained generalizable feed-forward Gaussian networks.
Figure 2: The composition of our training set and testing set. The training set consists of three RGB cameras (’front’, ’left-front’, and ’right-front’) mounted on our vehicle lane. We sample image sequences along trajectories that run parallel to the vehicle’s route for the test set. This encompasses four distinct test trajectories, each offset from the vehicle’s trajectory by 0 meters, 1 meter, 2 meters, and 4 meters, respectively. The sampling interval for the test set is five times the sampling interval used for the training set.
@@ 2.4 Datasets in Autonomous Driving
In autonomous driving training and benchmarking, there are many datasets available. For example, KITTI [22]. KITTI-360 [34], vKITTI [8], CityScapes [17], Mapillary [39], ApolloScape [25], Waymo Open Dataset [46]. nuScenes [9], Argoverse [10] and Argoverse 2 [54], BDD100K [63], OpenLane-V2 [52], LiDAR-CS [20], etc.. Those previous works have laid the groundwork for research and development in autonomous driving algorithms. A comprehensive survey of datasets related to autonomous driving refers to [36]. More recently, a newly proposed Hierarchical 3D-GS [27] showcases the scalability of 3D-GS followed by a short scene with parallel trajectories to train and render. However, none of these existing researches have specifically addressed the evaluation of novel view synthesis techniques tailored for autonomous driving simulation, particularly in terms of their ability to meet the high demands of cross-lane capability.
@@ 3 Dataset
Distinguishing our dataset from previous datasets (both real-world and synthetic) that typically capture only a single road trajectory, ours additionally captures multiple parallel trajectories. To this end, we need to generate cross-lane data in the created worlds; here, we utilize Carla [19], an autonomous driving simulator platform built on Unreal Engine.
@@ 3.1 Sensor Setup
To meet the needs of most NeRF and 3D-GS algorithms, the sensors used to capture data include three color cameras (i.e. ‘left-front’, ‘front’, ‘right-front’) and one 3D laser scanner. The spatial relationships between sensors and the vehicle are fixed as shown in Fig. 2. All three RGB cameras share identical intrinsic, lens parameters. Specifically, they have a resolution of $1920\times 1280$, a field of view (FOV) of $49.5^{\circ}\times 36.7^{\circ}$, a sensitivity of ISO 100, and a shutter speed of 5ms. The 3D laser scanner features 64 laser beams, a scanning range of 60 meters, and laser beam angles ranging from $-30^{\circ}$ to $30^{\circ}$ in the vertical direction.
@@ 3.2 Data Generation
Our simulation environment consists of six scenes (“Town01,” “Town02,” “Town03,” “Town04,” “Town05,” and “Town10”, all of them are under CC-BY License provided by [19]) with various weather conditions, such as sunshine and rain, closely resembling real-world settings. An overview of all the scenes and the trajectories is shown in Fig. 1. For example, the training set samples 150 times in one scene to capture the images of three cameras, LiDAR points, and the vehicle’s extrinsic information. The vehicle’s forward distance between the two sample points is 2 meters. For the evaluation process, in contrast to datasets captured in real-world scenarios, our cameras are configured with identical parameters, ensuring uniform image quality across all cameras. Therefore, we only test the novel-view-synthesis in the view of front camera, which includes four groups. These images are aligned parallel to the training set, with each group exhibiting a progressive deviation of 0m, 1m, 2m, and 4m along the y-axis in vehicle coordinate, as shown in Fig. 2. Additionally, a few novel-view-synthesis for AD methods like GaussianPro [16], MARS [57], and UC-NeRF [15] need annotated sky masks to split the scene into foreground-sky and model the color compensation separately. We employ a pre-trained SegFormer [58] to effectively infer semantic segmentation masks and extract sky masks from them.
@@ 4 Benchmark
Table 1: Results on our proposed dataset with the different offsets using front-only camera. $\uparrow$: higher is better, $\downarrow$: lower is better. The red, orange, and yellow colors respectively denote the best, the second best, and the third best results.
@@
Table 2: Results on our proposed dataset with the different offsets using left-front, front, right-front cameras. $\uparrow$: higher is better, $\downarrow$: lower is better. The red, orange, and yellow colors respectively denote the best, the second best, and the third best results.
Method w/o Offset Offset-1m Offset-2m Offset-4m
[…]
- NeRF-based
- Gaussian-based
DC-Gaussian [53] 30.23 0.912 0.271 26.74 0.883 0.315 25.53 0.872 0.329 23.53 0.860 0.357
4.1 Benchmarking Environment
To comprehensively evaluate the performance and computational efficiency of the assessment methods, we conducted a series of experiments using an NVIDIA Tesla V100 16GB GPU. We benchmarked the selected methods across five different cities that include various driving scenarios. Our findings are presented through both qualitative and quantitative analyses.
Figure 3: Visualization of the rendered images using DC-Gaussian and EmerNeRF with different offsets (i.e. 0m, 2m, and 4m) in two scenes. The discriminated areas are highlighted, and the areas with better results are marked as $\square$, while worse results are marked as $\square$.
4.2 Benchmarking Methods
MARS [57]: We inherit most of the parameter settings as the original MARS implementation. We employ the RAdam optimizer with $1\times 10^{-3}$ learning rate. Since our scenes are static without moving objects, we disable $\mathcal{L}_{sem}$, and the rest of the loss functions remain the same. We train the model with 50,000 steps.
UC-NeRF [15]: We inherit most parameter settings from the original UC-NeRF implementation. We employ the AdamW optimizer with $2.5\times 10^{-3}$ learning rate. The weight of sky loss is set to $2\times 10^{-3}$. We train the model with 40,000 steps.
3DGS [26]: We use the implementation of NerfStudio to evaluate our dataset. We employ the AdamW optimizer with $1\times 10^{-3}$ learning rate. For stability, we “warm up” the computation in lower resolution. Specifically, we start the optimization using 4 times smaller image resolution and then upsample twice after 500 and 1000 iterations. We train the model with 30,000 steps.
PVG [13]: We employ the Adam optimizer and maintain a similar learning rate for most parameters as the original PVG implementation. At the same time, we adjust the learning rate of the amplitude $\mathbf{A}$, opacity decaying $\mathbf{\beta}$ and opacity $\mathbf{O}$ to $3\times 10^{-5}$, $0.02$ and $0.005$ respectively. We train the model with 30,000 steps.
GaussianPro [16]: In alignment with the approach described in GaussianPro, our models are trained for 30,000 iterations across all scenes following GaussianPro’s training schedule and hyper-parameters. The interval step of the progressive propagation strategy is set to 50 where propagation is performed 3 times. The threshold $\sigma$ of the absolute relative difference is set to 0.8. We set $\beta=0.001$ and $\gamma=0.001$ for the planar loss.
4.3 Used Metrics
We adopt the evaluation criteria employed by the methods mentioned above, which comprise Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) as our evaluation metrics. Furthermore, we ensure transparency and clarity by describing the experimental framework employed to compare the methods.